OpenStudy (anonymous):

what is the sum of a-1/abc^3+3-b/abc^3

OpenStudy (anonymous):

lcd is abc^3

OpenStudy (anonymous):

\[\frac{ (a-1)(3-b) }{ abc^3 }\]

OpenStudy (phi):

when you add fractions with the same denominator, you add their tops

OpenStudy (anonymous):

a-4-b?

OpenStudy (anonymous):

\[\frac{ a-4-b }{ abc^3 }\]

OpenStudy (phi):

first write it down a-1+3-b now simplify

OpenStudy (phi):

you can also write it as a+ -1 + 3 + -b and switch the order a + -b + 3 + -1

OpenStudy (phi):

3+ -1 or 3-1 is 2

OpenStudy (anonymous):

\[\frac{ a-b+2 }{ abc^3 }\]

OpenStudy (phi):

yes

OpenStudy (anonymous):

ok ty